OpenStudy (anonymous):

simplify the fraction of 20/28

OpenStudy (anonymous):

2/14

OpenStudy (anonymous):

20/2 = 10 28/2 = 14 10/2 = 5 14/2 = 7 Simplified fraction: 5/7